# Complements of discriminants of real parabolic function singularities. II

### v2

@@ 4.2. The homological invariant of real and virtual components
Let $f_{0}$ be a function singularity of class $P_{8}^{1}$ or $P_{8}^{2}$ having the form (41). Let $D$ be a small neighborhood of the point $\{f_{0}\}$ in the space $\Xi_{1}$ or $\Xi_{2}$. Let $B$ be a sufficiently large (compared to $D$) closed ball in ${\mathbb{R}}^{3}$, and let $f_{\lambda}$, $\lambda\in D$, be a small perturbation of $f_{0}$. Define $T_{-}(\lambda)$ as a real number slightly lower than the minimum of 0 and the lowest real critical value of $f_{\lambda}$. Similarly, define $T_{+}$ as a number slightly higher than the maximum of 0 and the highest real critical value of $f_{\lambda}$. For any $\lambda\in D$, define three sets: $W_{\lambda,0}=f_{\lambda}^{-1}((-\infty,0])\cap B$, $W_{\lambda,-}=f_{\lambda}^{-1}((-\infty,T_{-}])\cap B,$ and $W_{\lambda,+}=f_{\lambda}^{-1}((-\infty,T_{+}])\cap B$. The sets $W_{\lambda,-}$ and $B\setminus W_{\lambda,+}$ form then a trivializable fiber bundle over $D$. These sets contain their intersections with $\partial B$ as deformation retracts. In the case $P_{8}^{1}$ these intersections are just the discs. In the case $P_{8}^{2}$ each of them homeomorphic to the disjoint union of a disc and an annulus. In particular, the reduced homology groups $\tilde{H}_{i}(W_{\lambda,-})$ and $\tilde{H}_{i}(B\setminus W_{\lambda,+})$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
@@
The homology index $\mbox{HI}(f_{\lambda})$ of a non-discriminant function $f_{\lambda}$, $\lambda\in D$ is the triple of Betti numbers $b_{0},b_{1},b_{2}$ of the relative homology group $H_{*}(W_{\lambda,0},W_{\lambda,-})$.
@@
(43) $(W_{\lambda,+},W_{\lambda,0},W_{\lambda,-}).$
[…]
@@ 4.3. The symmetry group and the relation between the real and virtual components
The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following exceptions: the frames of vanishing cycles in the 2-homology groups of zero-level sets of polynomials consist of eight elements, and the data set of virtual functions only includes the parities of the Morse indices of critical points, and not their integer values. We also have the following analog of Proposition 2
@@
1. Any virtual Morse function of the type $\Xi_{1}$ or $\Xi_{2}$ is associated with a real polynomial of the same type.
2. If two virtual Morse functions of type $P_{8}^{1}$ (respectively, $P_{8}^{2}$) belong to the same virtual component, then they are associated with two generic Morse polynomials of class $\Xi_{1}$ (respectively, $\Xi_{2}$) that are isotopic to each other within the set of Morse polynomials. $\Box$
@@
In particular, this group can act on the set of the connected components of the discriminant complement in the spaces $\Xi_{i}$. The transformations (44) preserve all topological characteristics of the functions. Thus, all connected components in the same orbit of this action are associated with the same virtual component of the formal graph of type $P_{8}^{1}$ or $P_{8}^{2}$.
@@
also acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. It preserves the invariant Ind and defines involutions on both sets of real and virtual components of these classes.
@@
If two generic polynomials, $f$ and $\tilde{f}$, of class $\Xi_{1}$ or $\Xi_{2}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of this class that contain $f$ and $\tilde{f}$ are mapped to each other by an element of the group $S(3)$.
@@
For any polynomial $f:{\mathbb{R}}^{3}\to{\mathbb{R}}$ of degree three with principal homogeneous part of the form (41), ${\bf T}(f)$ is the unique polynomial of the form (42) with the same principal homogeneous part, which is obtained from $f$ by a parallel translation in ${\mathbb{R}}^{3}$. Namely, it is the translation by the vector $\frac{1}{3}(a,b,c)$, where $a,b,$ and $c$ are the coefficients of $f$ at the monomials $x^{2}$, $y^{2}$, and $z^{2}$.
@@
and two other polynomials obtained from it by permuting the coordinates.
Proof. According to Lemma 35 of [22], there are only six polynomials of form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$. Three of these polynomials (mentioned in formula (26) of [22]) are locally equivalent to $\xi^{3}-\eta^{2}-\zeta^{2}+\frac{64}{3}$ at the $A_{2}$ point. The remaining three polynomials have the form (46), in particular belong to the same orbit of the group $S(3)$. $\Box$
@@
Proof. This follows from statement 2 of Lemma 44 of [22] (the coefficients $\varkappa$ and $c$ in formula (45) of [22] are fixed by our conditions on the critical values). $\Box$
Proof of Proposition 33. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 34 (respectively, 35) repeats the deduction of Proposition 8 from Lemma 9. $\Box$
@@ 4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
Each virtual component of type $P_{8}^{1}$ is associated with only one connected component of the space of non-discriminant polynomials of class $\Xi_{1}$. So, there are exactly seven such connected components.
Proof. The invariant Ind of the polynomials
@@
is equal to 0 for C<6C<-6 and $C>6$, to $1$ for $C\in(-6,-2)$ and $C\in(2,6)$, and to 2-2 for $C\in(-2,2)$. These polynomials are invariant under the action of the group $S(3)$ of permutations of coordinates. Therefore, all virtual components represented by these polynomials are associated with single real connected components of the set of non-discriminant polynomials.
Two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2. For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$. According to the previous paragraph, the associated virtual components are also different.
@@
have $\mbox{Ind}=-3$. They are related to each other by the involution (45). Hence, by the last statement of Proposition 36, they realize two different virtual components. Since these polynomials are invariant under the entire group $S(3)$, these virtual components are also associated with single real components.
@@
in some local coordinates, and a Morse critical point with Morse index $1$ and negative critical value.
Proof. 1. At each critical point of the polynomial (51), the sum of its three partial derivatives is equal to $3(x^{2}+y^{2}+z^{2}),$ so the origin is the only critical point of (51) in the real domain. In coordinates $l,p_{l},q_{l}$, this polynomial has the form
@@
The lowest quasihomogeneous part of the polynomial (54) in the coordinates $l,p_{l},$ and $q_{l}$ with weights $\deg l=3$ and $\deg p_{l}=\deg q_{l}=2$ is equal to $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}.$ Therefore, by the techniques of [2], §12, this polynomial has the normal form $Z^{2}+X^{3}+XY^{2}$ of $D_{4}^{+}$ class in some local coordinates.
2. Let $\alpha$ be the coefficient of the polynomial $\psi_{2}$ at the monomial $q_{l}^{2}$. Then the lowest quasihomogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$. The substitution $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2},P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial to the polynomial with the lowest quasihomogeneous part
@@
Again, using the techniques of §12 of [2], this function can be reduced to the standard form (53) of $A_{3}$ singularities via a local change of coordinates. If $\varepsilon$ is sufficiently small, then adding $\varepsilon p_{l}^{2}$ to (51) does not bring four non-real critical points of (51) back to the real domain. On the other hand, by the index considerations the obtained polynomial $F(l)$ has one Morse critical point in addition to the $A_{3}$ point at the origin. This Morse critical point depends continuously on $l$, therefore its Morse index is the same for all $l$. It is easy to calculate that for $l\equiv x$ it is equal to 1. Our critical point of class $A_{3}$ is topologically equivalent to a Morse critical point with Morse index 2 (and can be perturbed to such a point by an arbitrarily small change of the polynomial $F(l)$). The Morse-like complex of the polynomial $F(l)$ is acyclic, therefore the critical value of this $A_{3}$ type critical point should be higher than that of the only other critical point of Morse index 1. $\Box$
@@
The polynomial $\tilde{F}(l)$ is canonically determined by the linear form $l$ and depends continuously on it. The set of all polynomials $\tilde{F}(l),$ $l\in L$, is connected and invariant under the action of the group $S(3)$. Thus the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement. $\Box$
4.5. One-dimensional cohomology of discriminant complements of $P_{8}^{1}$ singularities
@@
The first cohomology group of the component of the discriminant complement of the $P_{8}^{1}$ singularities that consists of polynomials with $\mbox{Ind}=-1$ is non-trivial.
Proof. The zero-level set of the polynomials in this component is homeomorphic to the plane with a handle. Indeed, the topology of these sets remains constant along the component. For any polynomial $\tilde{F}(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to the plane) by passing through a Morse critical point of Morse index 1 once.
In particular, the first homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$. These homology groups form a local system over the component, in particular the fundamental group of the component acts on them by monodromy. For each $l\in L$, one basic homology class of the zero-level set of $\tilde{F}_{l}$ is defined by the intersection of this set and the plane $\ker l$. When $l\in L$ runs a half of the circle $L$, the plane $\ker l$ runs a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomial $\tilde{F}(l)$ runs a closed cycle in the discriminant complement. The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation. $\Box$
4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularity
@@
The invariant Card of a virtual component is the number of its vertices.
The involution (45) preserves this invariant.
@@
1. There are exactly 9174 distinct virtual functions of $P_{8}^{2}$ singularity class.
2. These virtual functions are distributed over exactly fifteen virtual components. The values of invariants Ind and Card of these components are shown in the second and third rows of Table 3.
3. All pairs of different virtual components of $P_{8}^{2}$ singularity with the same invariant Card value can be realized by generic polynomials connected by involution ((45)).
@@
Each of the ten virtual components of the formal graph of type $P_{8}^{2}$ that are mentioned in columns 1, 3, 4, 6, 7, 11, 12, 13, 14, and 15 of Table 3 is associated with only one real connected component of the discriminant complement.
@@
All of the polynomials given in the next formulas (55), (56), (57), and (58) are invariant under the $S(3)$ action. Thus, for any polynomial $f$ of this list and any noncritical value $C$ of $f$, the set-theoretical invariant of the polynomial fCf-C is a virtual component associated with only one real component of the discriminant complement. We will show that these virtual components, or those obtained from them via the involution (45), include all virtual components with Card equal to 258, 336, 1648, 262, and 94. The proof of Theorem 41 for the components with $\mbox{Card}=60$ is more implicit, see § 43.
Components with $\mbox{\rm Ind}=0$ and $\mbox{\rm Card}=1648$. For each polynomial $f$ of class $\Xi_{2}$, the polynomials $f+C$ and fCf-C with sufficiently large $C$ realize two different virtual components with $\mbox{Ind}=0$. Indeed, the homology index of one of these components is $(000)$, and the other’s is $(011)$. In particular, the $S(3)$-invariant polynomials
(55) $x^{3}+y^{3}+z^{3}-4xyz+C\qquad\mbox{and}\qquad x^{3}+y^{3}+z^{3}-4xyz-C$
represent these two components for any $C>0$. These polynomials are not Morse, but any their sufficiently small generic perturbations are associated with virtual functions with $\mbox{Card}=1648$.
Components with $\mbox{\rm Card}=262$. The polynomial
@@
has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2. Thus, it has invariant values $\mbox{Ind}=1$ and $\mbox{HI}=(00\bar{1})$. This polynomial is invariant under the involution (45). According to the last statement of Proposition 40, it cannot represent a component of the discriminant complement with $\mbox{Card}=94$. Therefore, it belongs to the component with $\mbox{Card}=262$.
@@
(57) $f\equiv x^{3}+y^{3}+z^{3}-6xyz-\frac{3}{2}(x^{2}+y^{2}+z^{2})$
has eight real Morse critical points: one local maximum with critical value 0, three points of Morse index $2$ with value $\frac{11}{49}$, three points of index $1$ with value $\frac{1}{2}$, and one more point of index $2$ with value $\frac{3}{2}$. Therefore, the index $\mbox{Ind}(f-C)$ is equal to 0 for $C<-\frac{3}{2}$, to 1 for $C\in\left(-\frac{3}{2},-\frac{1}{2}\right),$ to 2-2 for $C\in\left(-\frac{1}{2},-\frac{11}{49}\right),$ to $1$ for $C\in\left(-\frac{11}{49},0\right),$ and to 0 for $C>0$.
The group $H_{*}(W_{\lambda,+},W_{\lambda,0})$ corresponding to the polynomials fCf-C with $C\in\left(-\frac{11}{49},0\right)$ is isomorphic to $\mathbb{Z}$} in dimension three and is trivial in all other dimensions. Therefore, by the exact sequence of the triple (43), the homology index $\mbox{HI}(f+\frac{1}{7})$ is equal to $(0,1,2)$. The involution (45) moves this polynomial to that with $\mbox{HI}=(1,0,0)$. In particular, these two polynomials represent two different components, and the associated virtual components are also different. According to Table 3, these are the components with $\mbox{Card}=94$.
The homology index $\mbox{HI}(f-C)$ with $C\in\left(-\frac{3}{2},-\frac{1}{2}\right)$ is $(0,0,1),$ because the corresponding group $H_{*}(W_{\lambda,0},W_{\lambda,-})$ is generated by the class of one critical point of Morse index $2$. Therefore, it is again the component with $\mbox{Card}=262$.
[…]
Components with $\mbox{\rm Card}=258,$ $336,$ and $94$. The polynomial
(58) $f=x^{3}+y^{3}+z^{3}-12xyz-\frac{1}{2}(x^{2}+y^{2}+z^{2})$
has eight real Morse critical points: a local maximum with critical value 0, three points with Morse index 2 and critical value $\frac{11}{3969}$, one point more with Morse index 2 and critical value $\frac{1}{162}$, and three points with index 1 and critical value $\frac{1}{54}$. Thus, $\mbox{Ind}(f-C)$ is equal to 0 for $C<-\frac{1}{54}$, to 3-3 for $C\in\left(-\frac{1}{54},-\frac{1}{162}\right)$, to 2-2 for $C\in\left(-\frac{1}{162},-\frac{11}{3969}\right)$, to $1$ for $C\in\left(-\frac{11}{3969},0\right)$, and to 0 for $C>0$.
@@
Proof. By Proposition 40 of [22], the intersection matrix of vanishing cycles of the canonical basis for all generic polynomials of class $\Xi_{2}$ with eight real critical points, one of which is an extremum, is expressed by the Dynkin diagram shown in

### v3

@@ 4.2. The homological invariant of real and virtual components
Let $f$ be a polynomial in the class $\Xi_{1}$ or $\Xi_{2}$. Let $T_{-}$ (respectively, $T_{+}$) be an arbitrary negative (respectively, positive) constant which is lower (respectively, higher) than all critical values of $f$.
Consider three subsets in ${\mathbb{R}}^{3}$: $W_{0}=f^{-1}((-\infty,0])$, $W_{-}=f^{-1}((-\infty,T_{-}]),$ and $W_{+}=f^{-1}((-\infty,T_{+}])$. If $f\in\Xi_{1}$ (respectively, $f\in\Xi_{2}$) then the space $W_{-}$ is contractible (respectively, consists of two components, one homotopy equivalent to a point and the other to a circle). In particular, the reduced homology groups $\tilde{H}_{i}(W_{-})$ for $f\in\Xi_{2}$ are equal to $\mathbb{Z}$} for $i=0$ or 1 and are trivial for all other $i$.
@@
The homology index $\mbox{HI}(f)$ of a non-discriminant polynomial $f\in\Xi_{1}$ or $\Xi_{2}$ is the triple of Betti numbers $(b_{0},b_{1},b_{2})$ of the relative homology group $H_{*}(W_{0},W_{-})$.
@@
(43) $(W_{+},W_{0},W_{-}).$
The groups $H_{*}(W_{+},W_{-})$ participating in this exact sequence are the same for all $f$ from the same space $\Xi_{i}$, because the pairs of sets $(W_{+},W_{-})$ corresponding to different functions $f$ form a trivializable fiber bundle over these spaces. In the case $\Xi_{1}$ these groups are trivial. For $\Xi_{2}$, they are equal to $\mathbb{Z}$} for $i=1$ or $2$ and are trivial for all other $i$.
@@ 4.3. The symmetry group and the relation between the real and virtual components
The virtual functions, formal graphs, and virtual components of the $P_{8}^{*}$ singularities are defined almost identically to those of the $X_{9}$ singularities, see § 2.2, with only the following alterations: the frames of vanishing cycles in the groups $H_{2}(f^{-1}(0))$ for non-discriminant $f\in\Xi_{1}$ or $\Xi_{2}$ consist of eight elements, and the data set of virtual functions only includes the parities of the Morse indices of critical points, and not their integer values. We also have the following analog of Proposition 3.
@@
1. Any virtual function of the class $\Xi_{1}$ or $\Xi_{2}$ is associated with a real polynomial of the same type.
2. If two virtual functions of the class $\Xi_{1}$ ((respectively, $\Xi_{2})$ belong to the same virtual component, then they are associated with two generic polynomials of class $\Xi_{1}$ ((respectively, $\Xi_{2})$ that belong to the same connected component of the space of generic polynomials. $\Box$
@@
In particular, this group can act on the set of the connected components of the discriminant complements in the spaces $\Xi_{i}$. The transformations (44) preserve all topological characteristics of the functions. Therefore, all these connected components within the same orbit of this action are associated with the same virtual component of the formal graph of type $P_{8}^{1}$ or $P_{8}^{2}$.
@@
also acts on the spaces $\Xi_{1}$ and $\Xi_{2}$. It defines an involution on both sets of real and virtual components. This involution preserves the invariant Ind.
If $f$ and $\tilde{f}$ are two functions related by the involution (45), then the corresponding exact sequences (43) are related by a kind of Poincaré duality: the group $H_{*}(W_{+},W_{0})$ for the polynomial $f$ is naturally dual to $H_{*}(W_{0},W_{-})$ for $\tilde{f}$, and vice versa.
@@
If two generic polynomials, $f$ and $\tilde{f}$, of class $\Xi_{1}$ or $\Xi_{2}$ are associated with the same virtual function, then the connected components of the space of generic polynomials of this class containing $f$ and $\tilde{f}$ are mapped to each other by an element of the group $S(3)$. In particular, in this case the connected components of the discriminant complement containing these polynomials are also mapped to each other by an element of the group $S(3)$.
@@
For any polynomial $f:{\mathbb{R}}^{3}\to{\mathbb{R}}$ of degree three with principal homogeneous part of the form (41), ${\bf T}(f)$ is the unique polynomial of the form (42) with the same principal homogeneous part, which is obtained from $f$ by a parallel translation in ${\mathbb{R}}^{3}$. Specifically, it is the translation by the vector $\frac{1}{3}(a,b,c)$, where $a,b,$ and $c$ are the coefficients of $f$ at the monomials $x^{2}$, $y^{2}$, and $z^{2}$.
@@
and two other polynomials obtained from this one by permuting the coordinates.
Proof. According to Lemma 35 of [28], there are only six polynomials of form (42) that have a critical point of class $E_{6}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{64}{3}$. Three of these polynomials (mentioned in formula (26) of [28]) have signature (0,2) at the $A_{2}$ point. The remaining three polynomials have the form (46), in particular belong to the same orbit of the group $S(3)$. $\Box$
@@
Proof. This lemma follows almost immediately from statement 2 of Lemma 44 in [28] (the coefficients $\varkappa$ and $c$ in formula (45) of [28] are fixed by our conditions on the critical values). $\Box$
Proof of Proposition 34. The deduction of this proposition for the case $\Xi_{1}$ (respectively, $\Xi_{2}$) from Lemma 35 (respectively, 36) repeats the deduction of Proposition 9 from Lemma 10. $\Box$
@@ 4.4. Components of discriminant complements of $P_{8}^{1}$ singularities
Each virtual component of type $P_{8}^{1}$ is associated with one and only one connected component of the space of non-discriminant polynomials of class $\Xi_{1}$. In particular, there are exactly seven such connected components.
Proof. The invariant Ind of the polynomial
@@
is equal to 0 for C<6C<-6 and $C>6$, to $1$ for $C\in(-6,-2)$ and $C\in(2,6)$, and to 2-2 for $C\in(-2,2)$. These polynomials are invariant under the action of the group $S(3)$ of permutations of coordinates. According to Proposition 34, all virtual components represented by these polynomials are associated with single real connected components of the set of non-discriminant polynomials.
The two real components with $\mbox{Ind}=1$ corresponding to $C\in(-6,-2)$ and $C\in(2,6)$ are separated by the homology invariant of § 4.2. For the first component, this invariant is equal to $(100)$, and for the second it is equal to $(001)$. According to the previous paragraph, the associated virtual components are also different.
@@
have $\mbox{Ind}=-3$. They are related to each other by the involution (45). Therefore, according to the final statement of Proposition 37, they realize two distinct virtual components. Since these polynomials are invariant under the entire group $S(3)$, these virtual components are also associated with single real components. Thus, Theorem 38 is proved for all virtual components with Ind equal to 1,0,2,1,0,-2, and 3-3. Let us prove it for the only remaining virtual component with $\mbox{Ind}=-1$.
@@
in some local coordinates, and a Morse critical point with Morse index $1$ and a negative critical value.
Proof. 1. The sum of the three partial derivatives of the polynomial (51) is identically equal to $3(x^{2}+y^{2}+z^{2}),$ so the origin is the only critical point of (51) in the real domain. In coordinates $l,p_{l},q_{l}$, this polynomial has the form
@@
The lowest quasihomogeneous part of the polynomial (54) in the coordinates $l,p_{l},$ and $q_{l}$ with weights $\deg l=3$ and $\deg p_{l}=\deg q_{l}=2$ is equal to $l^{2}+p_{l}^{3}+p_{l}q_{l}^{2}.$ Therefore, by the techniques of [3], §12, this polynomial can be reduced to the normal form $\tilde{z}^{2}+\tilde{x}^{3}+\tilde{x}\tilde{y}^{2}$ of $D_{4}^{+}$ class in some local coordinates $\tilde{x},\tilde{y}$, and $\tilde{z}$.
2. Let $\alpha$ be the coefficient of the polynomial $\psi_{2}$ at the monomial $q_{l}^{2}$. Then the lowest quasihomogeneous part of the polynomial $F(l)$ with weights $\deg p_{l}=2,\deg q_{l}=1,$ and $\deg l=2$ is equal to $l^{2}-\varepsilon p_{l}^{2}+p_{l}q_{l}^{2}+\alpha lq_{l}^{2}$. Substituting $\tilde{l}=l-\frac{\alpha q_{l}^{2}}{2}$ and $P_{l}=p_{l}-\frac{q_{l}^{2}}{2\varepsilon}$ turns this polynomial into the polynomial with the lowest quasihomogeneous part
@@
Again, using the techniques of §12 of [3], this function can be reduced to the standard form (53) of $A_{3}$ singularities via a local change of coordinates. If $\varepsilon$ is sufficiently small, then adding $\varepsilon p_{l}^{2}$ to (51) does not return the non-real critical points of (51) to the real domain. By the index considerations, the obtained polynomial $F(l)$ has one real Morse critical point in addition to the $A_{3}$ point at the origin. This Morse critical point depends continuously on $l$, therefore its Morse index is the same for all $l\in L$. It is easy to calculate that for $l\equiv x-y$ it is equal to 1.
The restriction of the polynomial $F(l)$ to the line connecting the origin and this Morse point is a degree three polynomial with vanishing derivative at these points and the local maximum at the $A_{3}$ point. Therefore the critical value at its Morse critical point is negative. $\Box$
@@
The polynomial $\tilde{F}(l)$ is canonically determined by the linear form $l$ and depends continuously on it. The union of all polynomials $\tilde{F}(l),$ $l\in L$, is connected and invariant under the action of the group $S(3)$. Therefore, the virtual component with $\mbox{Ind}=-1$ is also associated with only one component of the discriminant complement. $\Box$
4.5. One-dimensional cohomology of the discriminant complement of $P_{8}^{1}$ singularities
@@
The first cohomology group of the connected component of the discriminant complement in $\Xi_{1}$ that consists of polynomials with $\mbox{Ind}=-1$ is non-trivial.
Proof. The zero-level sets of the polynomials in this component are homeomorphic to the plane with a handle. Indeed, the topology of these sets remains constant along the component. For any polynomial $\tilde{F}(l)$ from the previous proof this set is obtained from the very low level sets (homeomorphic to a plane) by passing through a Morse critical point of Morse index 1.
Thus, the one-dimensional homology groups of all zero-level sets of class $\Xi_{1}$ polynomials with $\mbox{Ind}=-1$ are isomorphic to ${\mathbb{Z}}^{2}$. These homology groups form a local system over the space of all such polynomials, i.e., over the connected component under consideration. In particular, the fundamental group of this component acts on these groups by monodromy. For each $l\in L$, one basic homology class of the zero-level set of the polynomial $\tilde{F}_{l}$ is defined by the intersection of this set and the plane $\ker l$. When $l\in L$ runs a half of the circle $L$, the plane $\ker l$ runs a closed cycle in ${\mathbb{R}P}^{2}$, and the polynomial $\tilde{F}(l)$ runs a closed cycle in the discriminant complement. The monodromy over the latter cycle brings our basic 1-cycle to itself with the opposite orientation. Thus, the half-circle $L/{\mathbb{Z}}_{2}$ defines non-trivial elements of the fundamental group and of the mod 2 homology group of the component. $\Box$
4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
@@
The invariant Card of a virtual component of the formal graph of an arbitrary singularity is the number of vertices in that component.
The involution (45) of virtual components of classes $\Xi_{1}$ and $\Xi_{2}$ preserves this invariant.
@@
1. There are exactly 9174 distinct virtual functions of class $\Xi_{2}$.
2. These virtual functions are distributed among fifteen virtual components. The values of invariants Ind and Card of these components are shown in the second and the third rows of Table 3.
3. All pairs of different virtual components of class $\Xi_{2}$ with the same Card value ((i.e., the pairs of components 3 and 4, 6 and 7, 9 and 10, 11 and 12, and 14 and 15 in Table 3)) can be realized by generic polynomials connected by involution ((45)).
@@
Each virtual component of the formal graph of type $P_{8}^{2}$ is associated with only one real connected component of the discriminant complement. In particular, there are exactly fifteen such connected components.
@@
All of the polynomials given in the formulas (55), (56), (57), and (58) below are invariant under the $S(3)$ action. Thus, if $f$ is any of these polynomials and $C$ is a noncritical value of $f$, then Proposition 34 implies that the virtual component associated with the polynomial fCf-C is associated with only one real component of the discriminant complement. We will show that all virtual components with Card equal to 258, 336, 1648, 262, or 94 can be realized in this way. Hence, the assertion of Theorem 42 is true for these components.
Components with $\mbox{\rm Card}=1648$. For each polynomial $f$ of class $\Xi_{2}$, the polynomials fCf-C with sufficiently large $|C|$ and different signs of $C$ realize two different virtual components with $\mbox{Ind}=0$. The homology indices of these components are equal to $(000)$ and $(011)$. For example let $f$ be the $S(3)$-invariant polynomial
(55) $x^{3}+y^{3}+z^{3}-4xyz.$
The polynomial fCf-C with arbitrary $C\neq 0$ then represents one or the other of these two components depending on the sign of $C$. These polynomials are not Morse, but all their sufficiently small generic perturbations have $\mbox{Card}=1648$.
Component with $\mbox{\rm Card}=262$. The polynomial
@@
has only two real Morse critical points: one with Morse index 2 and critical value 2-2, and the other with Morse index 1 and critical value 2. Thus, it has $\mbox{Ind}=1$ and homology index $(001)$. This polynomial is invariant under the involution (45). Therefore, according to the last statement of Proposition 41, the associated virtual component cannot have $\mbox{Card}=94$. Thus, this is the only remaining component with $\mbox{Ind}=1$, which has $\mbox{Card}=262$.
@@
(57) $f\equiv x^{3}+y^{3}+z^{3}-6xyz+3(x^{2}+y^{2}+z^{2})$
has eight real Morse critical points: one local minimum with critical value 0, three critical points with Morse index $1$ and critical value $\frac{88}{49}$, three critical points with Morse index $2$ and critical value $4$, and one point with Morse index $1$ and critical value $12$. Therefore, the invariant $\mbox{Ind}(f-C)$ is equal to 0 for $C<0$, to $1$ for $C\in\left(0,\frac{88}{49}\right),$ to 2-2 for $C\in\left(\frac{88}{49},4\right),$ to 1 for $C\in\left(4,12\right),$ and to 0 for $C>12$.
The group $H_{*}(W_{0},W_{-})$ corresponding to any polynomial fCf-C with $C\in\left(0,\frac{88}{49}\right)$ is isomorphic to $\mathbb{Z}$} in dimension 0 and is trivial in all other dimensions. Therefore, the homology index $\mbox{HI}(f-1)$ is $(100)$. The involution (45) moves this polynomial to that with $\mbox{HI}=(012)$. In particular, these two polynomials represent two different components. They are $S(3)$-invariant, therefore by Proposition 34, the associated virtual components are also different. According to Proposition 41 and Table 3, these are two components with $\mbox{Card}=94$.
According to elementary Morse theory, the homology index of each polynomial fCf-C with $C\in\left(\frac{88}{49},4\right)$ is equal to $(020)$ or $(130)$, while its image under the involution (45) has $\mbox{HI}=(031)$. Therefore, these two polynomials belong to different components with $\mbox{Ind}=-2$ and equal Card invariants. The only possibility for this is two components with $\mbox{Card}=336.$
Component with $\mbox{\rm Card}=258$. The polynomial
(58) $f=x^{3}+y^{3}+z^{3}-12xyz+3(x^{2}+y^{2}+z^{2})$
has eight real Morse critical points: one local minimum with critical value 0; three critical points with Morse index 1 and critical value $\frac{88}{147}$; one critical point with Morse index 1 and critical value $\frac{4}{3}$; and three critical points with Morse index 2 and critical value $4$. Thus, $\mbox{Ind}(f-C)$ is equal to 0 for $C<0$, to $1$ for $C\in\left(0,\frac{88}{147}\right)$, to 2-2 for $C\in\left(\frac{88}{147},\frac{4}{3}\right)$, to 3-3 for $C\in\left(\frac{4}{3},4\right)$, and to 0 for $C>4$.
@@
Proof. By Proposition 40 of [28], the intersection matrix of the canonical basis of vanishing cycles for any generic polynomial of class $\Xi_{2}$ with eight real critical points, one of which is an extremum, is expressed by the Coxeter-Dynkin graph shown in